At a certain temperature a saturated solution of Zn(OH)2 has a pH of 8. What is the value of Ksp at this temperature?
1
Expert's answer
2019-04-25T07:58:51-0400

pH = 8, C(H+) = 10-8 M.

C(OH-) = Kw/C(H+) = 10-14/10-8 M = 10-6 M.


                   x M     10-6 M

Zn(OH)2 ⇄ Zn2+ + 2OH-

                   1 M      2 M


C(Zn2+) = x = 10-6 / 2 M = 0.5 × 10-6 M.

Ksp = C(Zn2+) × [C(OH-)]2 = 0.5 × 10-6 × (10-6)2 = 0.5 × 10-18.